Project

General

Profile

Update 11 April 2016

News: Physics and Step IV
Added by Kyberd, Paul about 2 months ago

Dear All,

The reviews of last week, together with the presentations the week before at CM44, provide an excellent summary of the status of the project.

I’m sending this brief update only to note the increments of progress made over the past week and to invite you all to the next “MICE briefing” which will be held in R9 on Thursday the 14th April 2016. The purpose of this briefing is to celebrate the hard work that allowed such an excellent impression to be given in the recent reviews and to allow us to discuss informally the issues raised by the revised framework for the recovery of the downstream solenoid.

With best wishes …

Ken

MICE briefing:
• R9; 15:00 on Thursday the 14th April 2016

Status update:

Physics and analysis:
The status of the analysis is succinctly summarised in the presentations to the MPB. The principal advancement over the past week is from Francois Drielsma who has identified issues in the handling of the magnetic field maps in the Monte Carlo. The issues relate to the geometrical “placement” of the maps. When these issues are corrected, the data/Monte Carlo comparison plots for the trackers and the TOFs are much improved. Further tuning will continue, but I am optimistic that the work for the emittance-reconstruction paper can now proceed with confidence.

We shall need to be careful to plan work on the emittance measurement and the scattering measurement such that we can prepare the preliminary results in time for the summer conferences. A practical hard deadline is afforded by the ICHEP 2016 paper-submission deadline which has just been extended to 31May16. Chris Rogers will be organising the analysis meetings such that the steps necessary for results to gain “Preliminary” status are carried out in a timely manner.

Step IV:
Of course, the hardware status is carefully summarised in the RLSR/MPB presentations. Progress beyond that noted there includes:
• The LiH absorber has been extracted from the focus-coil and moved to safe storage. Electrical work on the FC itself continues in preparation for the cool-down of the module later this month, which will be followed by stand-alone commissioning;
• Preparations to separate the “demin” and “inhibited” water circuits is underway. A chiller, loaned to us by ISIS, will be used to ensure a minimum of disruption to the circuit that cools the compressors serving the upstream and downstream solenoids.
• The remedial work on the water system has been given priority as the risk to the compressors of continued us of demineralised water is too high. This has the necessary consequence that weekend running in April 2016 has been cancelled. The weekend data-taking would have been used to “top up” the empty-channel data sets for the LiH scattering measurement. These data sets appear to be reasonably well populated and it seems likely that they will be sufficient to allow the scattering distributions to be unfolded.
• The installation of components of the QD and QP systems for the spectrometer solenoids continues. The necessary modifications to Rack Room 2 are underway, the DCCTs have been installed and the framework that will support the external dump resistors is being assembled in R9.

Future meetings:
• Video conferences:
• 05May16
• 09Jun16
• 14Jul16
• Physics workshop:
• 28th April 2016 at RAL


Comments